# Computational considerations for the prediction of airfoil Stall Flutter

### v1

@@ 3.2 The structural dynamics model
The structural response of the airfoil is handled internally by the in–house RBD solver of MaPFlow, which is integrated within the numerical framework MaPFlow and is strongly coupled with the aerodynamic system to perform high–fidelity FSI computations. For the validation study presented here, the RBD solver is strictly constrained to model only the pitch DOF. The governing equation of motion for pitch rotation $\theta$ is expressed as a standard second–order system:
@@ 3.4 Sensitivity analysis
In this section, a comprehensive sensitivity analysis of the employed numerical parameters is presented, aiming at minimizing numerical uncertainties and ensuring the robustness and validity of the predicted results. The core objectives of this section involve the rigorous investigation of grid and temporal discretisation effects. This study focuses specifically on a highly challenging LAO case at V=24m/s airspeed from the dimitriadis2009bifurcation experimental campaign, as this regime imposes the most stringent numerical requirements.
@@ 3.4.1 Grid Dependency Analysis
The initial focus of the sensitivity analysis is the refinement of the computational grid. An O–type mesh is utilised, extending to a far–field radius of one hundred chord lengths ($100c$) (see Figure 6). A refined near–wake region is employed to properly resolve the emitted vortices, defined by a characteristic length of $\Delta w=5*10^{-2}c$. This region extends $5c$ in the chordwise direction ($0.5c$ upstream and $4.5c$ downstream) and $4c$ in the normal direction ($2c$ towards the pressure and $2c$ towards the suction side). The airfoil surface is discretised in $N_{c}=1148$ cells along its perimeter, varying from a minimum spacing of $\Delta X_{min}=10^{-4}c$ at the leading and trailing edges to a maximum of $\Delta X_{max}=2*10^{-3}c$ close to the maximum thickness region, with an increase ratio of $1.1$. A structured mesh zone expands radially from the airfoil wall to ensure the correct prediction of the boundary layer development. The height of the first cell is $h_{1}=10^{-5}c$ (thereby ensuring a normalized wall distance of $y^{+}\leq 1$) and the total number of cells in the radial direction is $35$ with an increase ratio of $1.1$. For efficiency, an unstructured meshing technique is employed for the remainder of the domain, with a growth rate of $1.1$, resulting in a total number of approximately $105000$ cells.
@@
#cells\#\>cells Amplitude $^{\circ}$] Period [sec]
$\approx 60k$ $21.31^{\circ}$ $(-9.4\%)$ $1.3175$ $(+19.4\%)$
$\approx 105k$ $23.54^{\circ}$ $(+0.1\%)$ $1.1423$ $(+3.5\%)$
$\approx 215k$ $23.36^{\circ}$ $(-0.7\%)$ $1.1535$ $(+4.5\%)$
$\approx 440k$ $23.52^{\circ}$ ()(-) $1.1037$ ()(-)
@@
A specific analysis on the minimum spacing, $\Delta X_{min}$, has been also conducted, highlighting its critical influence on aeroelastic predictions. The effect of $\Delta X_{min}$ has been previously overlooked in the respective literature, where grid dependency analyses are usually restricted solely to computing static polars watrin2012computational; yabili2012unsteady; poirel2011computational; abbas2021aeroelastic; li2024mechanism. Table 2 provides a clear overview of the effect of $\Delta X_{min}$ on the Amplitude and Period of the predicted LCO, clearly indicating that $\Delta X_{min}=10^{-4}c$ yields effectively in a grid–independent solution. On the other hand, when the coarse leading and trailing edge discretisation ($\Delta X_{min}=10^{-3}c$) is employed, Pitch Amplitude is overestimated by approximately $6.7\%$ and the LCO Period is underestimated by approximately $3.7\%$. Figure 7 further illustrates this by showing the time history predicted when different values of $\Delta X_{min}$ are employed, where it is shown that the size of $\Delta X_{min}$ clearly affects the Amplitude and Period of the predicted pitching motion. This sensitivity stems from the influence of the grid resolution, particularly at the leading edge, on the precise development and subsequent detachment dynamics of the Leading Edge Vortex (LEV). Crucially, the flow field and pressure distribution over the airfoil presented in Figure 8 are captured at the time instant denoted by the red circle in Figure 7. At this specific time, the differences observed in the aerodynamic field and thus the resulting aerodynamic moment are mainly attributed to the $\Delta X_{min}$ selection, as the airfoil exhibits similar pitch angle (see Figure 7(a)) and angular velocity (see Figure 7(b)) across the tested discretisations up to this point. When the leading edge discretisation is coarser, the LEV finds it challenging to detach, remaining closer to the airfoil (see Figure 8(a)). This proximity leads to higher local velocities and thus increased suction aft from the pitching axis (see Figure 8(b)). Ultimately, this leads to an increased excitation aerodynamic moment at this time instant that later results in higher pitch angles (see the different peak values predicted in 7 right after the red dot). Conversely, finer discretisation facilitates smoother LEV shedding and convection, moving the vortex farther from the airfoil and thereby generating an increased restoring aerodynamic moment that drives the airfoil to lower pitch peak values after the red dot time instant. These complex, non–trivial dependencies, visually supported by the flow field structures presented in Figure 8, demonstrate conclusively that grid dependency analyses on Stall Flutter instabilities must be extended beyond standard criteria, focusing specifically on how dynamic flow phenomena like vortex shedding are affected, rather than merely relying on static airfoil polar predictions.
@@
Table 2: Variation of Amplitude $^{\circ}$] and Period [sec] with decreasing Leading Edge and Trailing Edge spacing Xmin[m]Xmin\>[m], defined with respect to airfoil chord $c$. Reference Amplitude and Period correspond to minimum value $Xmin=10^{-}5c$. $Xmin=10^{-}4c$ yields in a grid–independent solution.
@@
Figure 7: Impact of minimum spacing ($\Delta X_{min}$) at the leading and trailing edges on pitch angle (a) and pitch rate (b) time history. All three simulations have a similar pitch angle and pitch rate at this instant. Consequently, the deviations in the predicted pitch peak after the red dot can be attributed to different aerodynamic excitation at this time instant.
@@ 3.4.2 Time–step Dependency Analysis
$0.01$ $25.80^{\circ}$ $(+9.6\%)$ $0.971$ $(-14.8\%)$
$0.005$ $23.54^{\circ}$ $(+0\%)$ $1.1423$ $(+0.1\%)$
$0.0025$ $23.54^{\circ}$ ()(-) $1.1412$ ()(-)
(a) Dynamically pitching NACA0012 (imposed motion) at $a_{0}=10^{\circ}$, $a_{1}=15^{\circ}$ and $k=0.1$, at $Re=135000$ and $Ma=0.103$. Experimental data from lee2004investigation and numerical predictions from liggett2012temporal.
(b) Predictions by MaPFlow for an aeroelastically pitching NACA0012 in symmetric LAO at $Re=540000$ and $Ma=0.079$, firstly observed in dimitriadis2009bifurcation experimental campaign.
@@ 4.1 Small Amplitude Oscillations
In this sub–section, the aeroelastic response of a system that results in SAO is examined. This phenomenon has been observed by poirel2008self in a low and constrained transitional Reynolds number range (from $45000$ to $130000$). In poirel2011computational it is shown that the flow separation close to the trailing edge plays a determinant role in the onset of the fluttering motion, while the subsequent presence of a LSB stabilizes this motion into a LCO. MaPFlow predictions, illustrated in Figure 10, confirm this mechanism, showing that the emergence of trailing edge vortices at time–stand "11" drives the system towards an instability onset, which is subsequently stabilized by the formation of larger detached vortices near the trailing edge (time–stands "22" and "33"). Hence, the employment of transition modelling is considered essential to correctly capture the initial trailing edge separation, consequently leading to the SAO regime. Neglecting this transition model in the numerical simulations results in the complete suppression of the LCO.
@@
Figure 10: Pitch response time history, vorticity contours, friction and pressure coefficient distribution for a SAO case at a Reynolds number of $Re=1.17*10^{5}$, firstly observed in poirel2008self experimental campaign. Three critical time instants have been makred in the time history plot, namely: "11" where the emergence of trailing edge vortices leads to an instability onset that drives the system to oscillations, "22 and "33" where the subsequent formation of larger detached vortices stabilizes the system to LCO. The corresponding friction and pressure coefficient distributions imprint the flow dynamics associated with the vortices formation. Red dots in friction coefficient plots indicate the laminar separation positions where LSBs form.
@@
[…]
@@ 4.2 Large Amplitude Oscillations
The structural damping in the experimental setup (primarily due to friction in the pitch springs) was found to exhibit highly non–linear characteristics. Experimental observations, specifically mentioned in Section , indicated that the critical pitch damping coefficient $\zeta$ demonstrated pronounced non–linearity primarily near small pitch angles (around the zero position) and transitioned to a linear increase at higher pitch angles. This behavior necessitates the implementation of a non–linear damping model, typically structured as a function of the pitch angle $\zeta(\theta)$, as describe by Equation (9) and illustrated in Figure 12. Although the damping was noted to increase linearly at larger pitch angles, the component representing this linear increase has been entirely omitted in the final model. This simplification is justified by the fact that the corresponding angular velocity $\dot{\theta}$ is typically small at large angles, meaning that the contribution of this linear term to the overall off–wind response is insignificant. Omitting this term consequently reduces the difficulty of tuning the model by removing an additional unknown variable. Figure 13 presents the successful validation of the chosen non–linear damping function of Equation (9) against experimental free–decay tests. The comparison confirms that this model accurately captures the dynamic response of the system in the absence of significant aerodynamic excitation, illustrating the characteristic rapid decay of oscillations observed experimentally.
@@
Figure 12: Non–linear critical damping ratio $\zeta$ with respect to pitch angle $\theta$ (in degrees) employed in LAO simulations of dimitriadis2009bifurcation cases.
@@
Figure 14 illustrates the transition from static stability to LCO for increasing values airflow velocity. When the aeroelastic system is exposed to low airflow velocity, such as V=18m/sV=18\;m/s, the airfoil deflects to its respective static equilibrium angle by following the typical response of a $2nd$ order dynamic system. Although the simulations start with the airfoil at $0^{\circ}$ pitch angle, where negligible aerodynamic lift or moment is theoretically expected, even a minor asymmetry inherent in the computational grid can induce an initial asymmetric loading over the airfoil. Consequently, the airfoil deflects to a non–zero pitch angle, which subsequently establishes a non–zero lift calculation. Provided that the distance between the pitching axis and the quarter–chord is sufficiently large, and considering that the angular velocity of the resulting response is small, the aerodynamic moment around the pitching axis will overcome the restoring damping moment, causing the airfoil to deflect to progressively larger pitch angles, dynamically heading to the static equilibrium pitch angle. (In order to accelerate this initial phase of the simulations, the initial pitch angle of the airfoils analysed herein is set to $0.2^{\circ}$ nose–up.) As velocity increases to V=18.4m/sV=18.4\;m/s, the balance between aerodynamic energy input and system damping becomes critical, resulting in marginal stabilisation to the static equilibrium pitch angle. A further increase to V=18.6m/sV=18.6\;m/s leads to the observed numerical critical velocity, where self–starting and sustained oscillations are predicted. At this critical velocity, the simulation initially exhibits a bifurcation from the symmetric to the asymmetric aeroelastic mode, subsequently resulting in small amplitude asymmetric oscillations around the static equilibrium angle. For velocities exceeding $V_{c,num}=18.6\;m/s$, symmetric oscillations of significantly larger amplitude around the zero mean pitch angle are predicted. The phase plane plot presented in Figure 15 clearly depicts these trajectories, confirming the shift from stable static equilibrium (V<18.6m/sV<18.6\;m/s) to stable motion of asymmetric SAO (V=18.6m/sV=18.6\;m/s) or symmetric LAO (V>18.6m/sV>18.6\;m/s).
@@
Firstly, the numerical predictions clearly over–predict the critical velocity ($V_{c,num}=18.6\;m/s$), which is measured at $V_{c,exp}\approx 13\;m/s$ (see Figure 16(a)). Furthermore, the predicted pitch amplitude at the onset ($A_{c,num}\approx 14.3^{\circ}$) is substantially lower than the measured amplitude at the experimental critical speed ($A_{c,exp}\approx 34.1^{\circ}$). This under–prediction continues systematically across the measured velocity range, where the predicted plateau amplitude ($\approx 29^{\circ}$) remains approximately $13^{\circ}$ lower than the measured plateau ($\approx 42^{\circ}$). This amplitude discrepancy explains the larger oscillation period observed in Figure 16(b) and is directly linked to the lower aerodynamic excitation peaks ($C_{M}$ and $C_{L}$) predicted in the numerical simulations, as illustrated in Figures 16(c) and 16(d) respectively. It needs to be highlighted that, following the approach of li2007experimental, high frequency peaks are neglected in the process of identifying the maximum and minimum values of $C_{M}$ and $C_{L}$. Despite these differences in magnitude, the simulation successfully captures the overall qualitative trend of increasing amplitude followed by a plateau, and the corresponding reduction in period with increasing airflow velocity.
Furthermore, the sectional pitching moment and lift hysteresis curves (Figures 17 and 18) demonstrate strong qualitative agreement with the experimental data. The same outcome can be made also for Figure 19, where only the asymmetric part of the loop has been measured. Specifically, MaPFlow predicts quite well the minimum values of moment, along with the folding of the curve when the airfoil approaches the minimum pitch value. The disagreement in the respective maximum values of the pitching moment at positive pitch angles originates in the asymmetric shape of the experimental wing. Furthermore, the almost constant values of moment during the upstroke motion at negative pitch angles and the downstroke motion at positive angles is also successfully predicted. The high frequency ripples detected in the predicted moment around these almost constant values are not evident in measurements and are related to the break down of the large LEVs shown in time instants "44" and "88" of Figure 20, to the smaller structures shown in time instants "55" and "99". MaPFlow is also able to predict the extension of the linear phase in post–static–stall angles of attack, by successfully resolving the delay of the boundary layer separation from the airfoil surface during upstroke at positive pitch angles (see time instant "11" in Figure 20) and downstroke at negative angles (see time instant "66" in Figure 20). The closed loops predicted around $0^{\circ}$ pitch angle (they are more easily identified in moment curves) are also evident in measurements, but slightly shifted due to the experimental wing asymmetry. For the same reason, the measured slope of the linear part of the lift curves at positive pitch angles is slightly lower than that of the negative angles.
@@
Figure 16: Comparison of computationally predicted LCO characteristics against experimental measurements from dimitriadis2009bifurcation. Figures show the variation with airflow velocity of Pitch Amplitude (a), Period (b), maximum Pitching Moment Coefficient (c) and maximum Lift Coefficient (d). Systematic discrepancies include the over–prediction of the critical velocity ($V_{crit}$) and the under–prediction of the LCO amplitudes across the velocity range, along with the respective larger oscillation Period.

### v2

@@ 3.2 The structural dynamics model
The structural response of the airfoil is handled internally by an in–house RBD solver, which is integrated within the numerical framework MaPFlow and is strongly coupled with the aerodynamic system to perform high–fidelity FSI computations. For the validation study presented here, the RBD solver is strictly constrained to model only the pitch DOF. The governing equation of motion for pitch rotation $\theta$ is expressed as a standard second–order system:
@@ 3.4 Sensitivity analysis
In this section, a comprehensive sensitivity analysis of the employed numerical parameters is presented, aiming at deminishing numerical uncertainties and ensuring the robustness and validity of the predicted results. The core objectives of this section involve the rigorous investigation of grid and temporal discretisation effects. This study focuses specifically on a highly challenging LAO case at V=24m/sV=24\;m/s airspeed from the Dimitriadis and Li (2009) experimental campaign, as this regime imposes the most rigorous numerical requirements.
@@ 3.4.1 Grid Dependency Analysis
The initial focus of the sensitivity analysis is the refinement of the computational grid. An O–type mesh is utilised, extending to a far–field radius of one hundred chord lengths ($100c$) (see Figure 6). A refined near–wake region is employed to properly resolve the emitted vortices, defined by a characteristic length of $\Delta w=5*10^{-2}c$. This region extends $5c$ in the chordwise direction ($0.5c$ upstream and $4.5c$ downstream) and $4c$ in the normal direction ($2c$ towards the pressure and $2c$ towards the suction side). For what is considered the medium grid, the airfoil surface is discretised in $N_{c}=1148$ cells along its perimeter, varying from a minimum spacing of $\Delta X_{min}=10^{-4}c$ at the leading and trailing edges to a maximum of $\Delta X_{max}=2*10^{-3}c$ close to the maximum thickness region, with an increase ratio of $1.1$. A structured mesh zone expands radially from the airfoil wall to ensure the correct prediction of the boundary layer development. The height of the first cell is $h_{1}=10^{-5}c$ (thereby ensuring a normalized wall distance of $y^{+}\leq 1$) and the total number of cells in the radial direction is $35$ with a growth rate of $1.1$. For efficiency, an unstructured meshing technique is employed for the remainder of the domain, with a growth rate of $1.1$, resulting in a total number of approximately $105000$ cells.
@@
mesh type #cells\#\>cells Amplitude $^{\circ}$] Period [sec]
Coarse $\approx 60k$ $21.31^{\circ}$ $(-9.4\%)$ $1.3175$ $(+19.4\%)$
Medium $\approx 105k$ $23.54^{\circ}$ $(+0.1\%)$ $1.1423$ $(+3.5\%)$
Fine $\approx 215k$ $23.36^{\circ}$ $(-0.7\%)$ $1.1535$ $(+4.5\%)$
Extra Fine $\approx 440k$ $23.52^{\circ}$ ()(-) $1.1037$ ()(-)
@@
A specific analysis on the minimum spacing, $\Delta X_{min}$, has been also conducted, highlighting its critical influence on aeroelastic predictions. Table 2 provides a clear overview of the effect of $\Delta X_{min}$ on the Amplitude and Period of the predicted LCO, clearly indicating that $\Delta X_{min}=10^{-4}c$ yields effectively in a grid–independent solution. On the other hand, when the coarse leading and trailing edge discretisation ($\Delta X_{min}=10^{-3}c$) is employed, Pitch Amplitude is overestimated by approximately $6.7\%$ and the LCO Period is underestimated by approximately $3.7\%$. Figure 7 further illustrates this by showing the time history predicted when different values of $\Delta X_{min}$ are employed, where it is shown that the size of $\Delta X_{min}$ clearly affects the Amplitude and Period of the predicted pitching motion. This sensitivity stems from the influence of the grid resolution, particularly at the leading edge, on the precise development and subsequent detachment dynamics of the Leading Edge Vortex (LEV). Crucially, the flow field and pressure distribution over the airfoil presented in Figure 8 are captured at the time instant denoted by the red circle in Figure 7. At this specific time, the differences observed in the aerodynamic field and thus the resulting aerodynamic moment are mainly attributed to the $\Delta X_{min}$ selection, as the airfoil exhibits similar pitch angle (see Figure 7(a)) and angular velocity (see Figure 7(b)) across the tested discretisations up to this point (see Table 3). When the leading edge discretisation is coarser, the LEV finds it challenging to detach, remaining closer to the airfoil (see Figure 8(a)). This proximity leads to higher local velocities and thus increased suction aft from the pitching axis (see Figure 8(b)). Ultimately, this leads to an increased excitation aerodynamic moment at this time instant that later results in higher pitch angles (see the different peak values predicted in 7 right after the red dot). Conversely, finer discretisation facilitates smoother LEV shedding and convection, moving the vortex farther from the airfoil and thereby generating an increased restoring aerodynamic moment that drives the airfoil to lower pitch peak values after the red dot time instant. A similar qualitative outcome can be made for the trailing edge spacing. In fine TE discretisation, the TE shear layer has rolled up into a discrete detached vortex and a second one is already formed. For the crude dicretization the detachment of the TE vortex is delayed. These complex, non–trivial dependencies, visually supported by the flow field structures presented in Figure 8, demonstrate conclusively that grid dependency analyses on Stall Flutter instabilities must be extended beyond standard criteria, focusing specifically on how dynamic flow phenomena like vortex shedding are affected, rather than merely relying on static airfoil polar predictions.
@@
Table 2: Variation of Amplitude $^{\circ}$] and Period [sec] with decreasing Leading Edge and Trailing Edge spacing Xmin[m]Xmin\>[m]. Reference Amplitude and Period correspond to minimum value $Xmin=10^{-}5c$. $Xmin=10^{-}4c$ yields in a grid–independent solution.
@@
Table 3: Pitch angle and pitch rate values at the specific time instant denoted by the red circle in Figure 7 for different values of minimum spacing ($\Delta X_{min}$) at the leading and trailing edges. All three simulations have a similar pitch angle and pitch rate at this instant.
Xmin[]Xmin\>[-] Pitch angle $^{\circ}$] Pitch rate $[rad/s]$
$10^{-}3c$ $20.98$ $1.50$
$10^{-}4c$ $20.38$ $1.40$
$10^{-}5c$ $20.82$ $1.31$
@@
Figure 7: Impact of minimum spacing ($\Delta X_{min}$) at the leading and trailing edges on pitch angle (a) and pitch rate (b) time history. All three simulations have a similar pitch angle and pitch rate (see Table 3) at the specific time instant denoted by the red circle. Consequently, the deviations in the predicted pitch peak after the red dot can be attributed to different aerodynamic excitation at this time instant.
@@ 3.4.2 Time–step Dependency Analysis
$0.01$ $25.80$ $(+9.6\%)$ $0.971$ $(-14.8\%)$
$0.005$ $23.5$ $(+0\%)$ $1.1423$ $(+0.1\%)$
$0.0025$ $23.54$ ()(-) $1.1412$ ()(-)
(a) Dynamically pitching NACA0012 (imposed motion) at $a_{0}=10^{\circ}$, $a_{1}=15^{\circ}$ and $k=0.1$, at $Re=135000$ and $Ma=0.103$. Experimental data from Lee and Gerontakos (2004) and numerical predictions from Liggett and Smith (2012).
(b) Predictions by MaPFlow for an aeroelastically pitching NACA0012 in symmetric LAO at $Re=540000$ and $Ma=0.079$, firstly observed in Dimitriadis and Li (2009) experimental campaign.
@@ 4.1 Small Amplitude Oscillations
In this sub–section, the aeroelastic response of a system that results in SAO is examined. This phenomenon has been observed in a low and constrained transitional Reynolds number range (from $45000$ to $130000$). In Poirel et al. (2008) it is stated that the flow separation close to the trailing edge plays a determinant role in the onset of the fluttering motion, while the subsequent presence of a LSB stabilizes this motion into a LCO. Nevertheless, CFD predictions in Poirel et al. (2011) indicate a Kelvin–Helmholtz instability occurring in the shear layer close to the trailing edge as the onset mechanism to the fluttering motion and the resulting von Kármán vortex shedding as the LCO stabilisation factor. MaPFlow predictions, illustrated in Figure 10, confirm this mechanism, showing that the emergence of trailing edge shedding at time–stand "11" drives the system towards an instability onset, which is subsequently stabilized by the formation of larger detached vortices near the trailing edge (time–stands "22" and "33") that bring about the large unsteady fluctuations on the aft part of the airfoil shown in $C_{f}$ and $C_{p}$ diagrams. The presence of a LSB is also verified by the present study simulations in the positions where the wall shear stress vanishes (indicated by red dots). However, as shown in the Q–criterion contours displayed in Figure 11, a von Kármán vortex shedding is the sustaining factor of the fluttering motion that induces upstream–propagating instabilities in both sides of the airfoil. The employment of transition modelling is considered essential to correctly capture the initial trailing edge separation, consequently leading to the SAO regime. Neglecting transition in the numerical simulations results in the complete suppression of the LCO.
Figure 10: Pitch response time history, vorticity contours, friction and pressure coefficient distribution for a SAO case at a Reynolds number of $Re=1.17*10^{5}$, firstly observed in Poirel et al. (2008) experimental campaign. Three critical time instants have been makred in the time history plot, namely: "11" where the emergence of trailing edge vortices leads to an instability onset that drives the system to oscillations, "22 and "33" where the subsequent formation of larger detached vortices stabilizes the system to LCO. The corresponding friction and pressure coefficient distributions imprint the flow dynamics associated with the vortices formation. Red dots in friction coefficient plots indicate the laminar separation positions where LSBs form.
Figure 11: Pitch response time history, Q–criterion contours, friction and pressure coefficient distribution for a SAO case at a Reynolds number of $Re=1.17*10^{5}$, firstly observed in Poirel et al. (2008) experimental campaign. The presence of a LSB is computationally verified in $C_{f}$ plots, at the positions where the $C_{f}$ values drop to zero. Nevertheless, Q–criterion contours indicate a von Kármán vortex shedding as the LCO sustaining factor, forming upstream–propagating instabilities.
@@
Figure 12: Pitch amplitude (left) and frequency (right) variation with respect to Reynolds number. Experimental measurements from Poirel et al. (2008) and predictions with different turbulence modelling parameters by other CFD tools from Poirel et al. (2011). The Low Reynolds Correction and the Transitional models can fairly capture the increase in the amplitude at low Reynolds numbers ($Re\leq 85400$) and the following decrease as the Reynolds number further increases, whereas other predictions show a significant deviation from the measured trend. The predicted amplitudes are slightly larger than the measured ones in the latter Reynolds range for both computational options, which accounts for the corresponding lower frequencies predicted by the Transitional model for the same Reynolds range. Nevertheless, the trend of the increasing frequency along with the Reynolds number is generally predicted with accuracy.
@@ 4.2 Large Amplitude Oscillations
The structural damping in the experimental setup (primarily due to friction in the pitch springs) was found to exhibit highly non–linear characteristics. Experimental observations, specifically mentioned in Section 2.2, indicated that the critical pitch damping coefficient $\zeta$ demonstrated pronounced non–linearity primarily near small pitch angles (around the zero position) and transitioned to a linear increase at higher pitch angles. This behavior necessitates the implementation of a non–linear damping model, typically structured as a function of the pitch angle $\zeta(\theta)$, as describe by Equation (9) and illustrated in Figure 13.
@@
Figure 13: Non–linear critical damping ratio $\zeta$ with respect to pitch angle $\theta$ (in degrees) employed in LAO simulations of Dimitriadis and Li (2009) cases.
Although the damping was noted to increase linearly at larger pitch angles, the component representing this linear increase has been entirely omitted in the final model. This simplification is justified by the fact that the corresponding angular velocity $\dot{\theta}$ is typically small at large angles, meaning that the contribution of this linear term to the overall off–wind response is insignificant. Omitting this term consequently reduces the difficulty of tuning the model by removing an additional unknown variable. Figure 15 presents the successful validation of the chosen non–linear damping function of Equation (9) against experimental free–decay tests Do these additional comments suffice?]performed at an airspeed of 0, $2.5$ and 3.1m/s3.1\;m/s. The complete aeroelastic system has been analysed in these simulations. Aeroodynamics has been properly accounted for, even though the airflow velocity is very low, meaning that the only significant flow component comes from the pitching motion of the wing. Deflection from the initial resting position has been accomplished by applying an impulse, which can be modelled either as an initial non–zero pitch angle (see Figures 14(a) and 14(b)) or a sudden moment around the pitching axis (see Figure 15(a)). The comparison confirms that this model accurately captures the dynamic response of the system in the absence of significant aerodynamic excitation, illustrating the characteristic rapid decay of oscillations observed experimentally.
@@
Figure 16 and Figure 17 illustrate the transition from static stability to LCO for increasing values airflow velocity. When the aeroelastic system is exposed to low airflow velocity, such as V=18m/sV=18\;m/s, the airfoil deflects to its respective static equilibrium angle by following the typical response of a $2nd$ order dynamic system. Although the simulations start with the airfoil at $0^{\circ}$ pitch angle, where negligible aerodynamic lift or moment is theoretically expected, even a minor asymmetry inherent in the computational grid can induce an initial asymmetric loading over the airfoil. Consequently, the airfoil deflects to a near–zero pitch angle, which subsequently establishes a non–zero lift calculation. Provided that the distance between the pitching axis and the quarter–chord is sufficiently large, and considering that the angular velocity of the resulting response is small, the aerodynamic moment around the pitching axis will overcome the restoring damping moment, causing the airfoil to deflect to progressively larger pitch angles, dynamically heading to the static equilibrium pitch angle. (In order to accelerate this initial phase of the simulations, the initial pitch angle of the airfoils analysed herein is set to $0.2^{\circ}$ nose–up.) As velocity increases to V=18.4m/sV=18.4\;m/s, the balance between aerodynamic energy input and system damping becomes critical, resulting in marginal stabilisation to the static equilibrium pitch angle. A further increase to V=18.6m/sV=18.6\;m/s leads to the observed numerical critical velocity, where self–starting and sustained oscillations are predicted. At this critical velocity, the simulation initially exhibits a bifurcation from the symmetric to the asymmetric aeroelastic mode, subsequently resulting in small amplitude asymmetric oscillations around the static equilibrium angle. For velocities exceeding $V_{c,num}=18.6\;m/s$, symmetric oscillations of significantly larger amplitude around the zero mean pitch angle are predicted. The phase plane plot presented in Figure 17 clearly depicts these trajectories, confirming the shift from stable static equilibrium (V<18.6m/sV<18.6\;m/s) to stable motion of asymmetric SAO (V=18.6m/sV=18.6\;m/s) or symmetric LAO (V>18.6m/sV>18.6\;m/s).
@@
Figure 18: Comparison of computationally predicted LCO characteristics against experimental measurements from Dimitriadis and Li (2009). Figures show the variation with airflow velocity of Pitch Amplitude (a), Period (b), maximum Pitching Moment Coefficient (c) and maximum Lift Coefficient (d). Systematic discrepancies include the over–prediction of the critical velocity ($V_{crit}$) and the under–prediction of the LCO amplitudes across the velocity range, along with the respective larger oscillation Period.
Furthermore, the sectional pitching moment and lift hysteresis curves (Figures 19 and 20) demonstrate strong qualitative agreement with the experimental data. The same outcome can be made also for Figure 21, where only the asymmetric part of the loop has been measured. Specifically, MaPFlow predicts quite well the minimum values of moment, along with the folding of the curve when the airfoil approaches the minimum pitch value. The disagreement in the respective maximum values of the pitching moment at positive pitch angles originates in the asymmetric shape of the experimental wing. Furthermore, the almost constant values of moment during the upstroke motion at negative pitch angles and the downstroke motion at positive angles is also successfully predicted. The high frequency ripples detected in the predicted moment around these almost constant values are not evident in measurements and are related to the break down of the large LEVs shown in time instants "44" and "88" to the smaller structures shown in time instants "55" and "99" of Figure 22, where the dynamic stall mechanism of the flow is illustrated with the help of Vorticity contours. The delay of the boundary layer detachment that causes the extension of the linear part of the aerodynamic loads and the formation of the strong LEV that brings about their sharp drop are fairly resolved by the aeroelastic tool. MaPFlow is also able to predict the extension of the linear phase in post–static–stall angles of attack, by successfully resolving the delay of the boundary layer separation from the airfoil surface during upstroke at positive pitch angles (see time instant "11" in Figure 22) and downstroke at negative angles (see time instant "66" in Figure 22). The closed loops predicted around $0^{\circ}$ pitch angle (they are more easily identified in moment curves) are also evident in measurements, but slightly shifted due to the experimental wing asymmetry. For the same reason, the measured slope of the linear part of the lift curves at positive pitch angles is slightly lower than that of the negative angles.